A cylindrical container. How to find the volume of a container if you know the Surface area is 105 in squared and the radius is 3 in.
Answer by mananth(16946)

105 in squared and the radius is 3 in.
The curved surface area of a cylinder is 2*pi*r*h
r= radius, h= height of cylinder
105=2*pi*3*3*h
105/2*3*pi=h
Volume of cylinder = pi*r^2*h
V= pi*3*3*(105)/6*pi
V= 9*105/6
V=157.5 cu.in